### v2

# DrugAgent: Explainable Drug Repurposing Agent with Large Language Model-based Reasoning

Abstract
Drug repurposing, the identification of new therapeutic applications for existing drugs, offers a promising approach to accelerate drug development. This paper introduces a novel multi-agent framework that enhances the drug repurposing process by leveraging state-of-the-art machine learning techniques and integrating diverse knowledge sources. Our framework consists of three specialized agents: an AI Agent that develops robust drug-target interaction (DTI) models; a Knowledge Graph Agent that extracts DTIs from established databases such as DGIdb, DrugBank, CTD, and STITCH; and a Search Agent that engages with biomedical literature to annotate and verify computational predictions. By synthesizing outputs from these agents, our system harnesses a wide array of data sources to propose viable repurposing candidates. Preliminary results demonstrate the framework’s potential to not only predict drug-disease interactions but also to reduce the time and costs associated with traditional drug discovery methods. Our approach outperforms existing methods in predicting drug repurposing potential while providing interpretable results. This work highlights the scalability of multi-agent systems in biomedical research and their capacity to drive innovation in drug repurposing. The framework presented in this paper offers a promising pathway towards more efficient and cost-effective drug discovery processes, paving the way for accelerated therapeutic developments. Code is available https://anonymous.4open.science/r/DrugRepurposeLLM-E0B5/.
1 Introduction
Drug repurposing, also known as drug repositioning, is a strategy for identifying new uses outside the scope of the original medical indication of existing drugs. This approach has gained significant attention due to its potential to reduce the time, cost, and risk associated with drug development (Pushpakom et al., 2019; Chen et al., 2024b; Fu et al., 2022b). With the growing complexity of biological data and the increasing availability of diverse biomedical information sources (Chen et al., 2024a; Wu et al., 2022b), there is a pressing need for innovative computational strategies that can integrate and analyze these vast datasets (Huang et al., 2021; Lu, 2018).
Recent advances in artificial intelligence (AI), particularly in machine learning and knowledge graphs (Gyori et al., 2017), have shown great promise in addressing these challenges (Vamathevan et al., 2019). However, the integration of heterogeneous data sources and the effective interpretation of their interrelations remain significant hurdles. To overcome these obstacles, we propose a multi-agent system framework, where each agent specializes in a specific aspect of the drug repurposing process. This approach not only enhances the scalability and flexibility of the analysis but also allows for more robust and accurate predictions.
Our framework includes three primary agents: the AI agent, the Knowledge Graph (KG) Agent, and the Search Agent. The AI agent employs the DeepPurpose package (Huang et al., 2020) to develop and refine drug-target interaction (DTI) models. The knowledge graph Agent utilizes the drug-gene interaction database (DGIdb) (Cannon et al., 2024), DrugBank (Knox et al., 2024), Comparative Toxicogenomics Database (CTD) (Davis et al., 2023), and Search Tool for Interactions of Chemicals (STITCH) (Kuhn et al., 2007) to extract and synthesize information on DTIs. The search Agent engages with biomedical literature, leveraging large language models for automated data labeling and validation, a process resource-intensive and costly.
This paper details the design, implementation, and initial outcomes of this multi-agent system. By demonstrating the system’s ability to leverage AI for effective drug repurposing, we aim to contribute to the broader field of computational drug discovery, offering a scalable model that can be adapted and extended to other areas of biomedical research.
Unlike previous approaches that rely on single models or data sources, our multi-agent system leverages diverse perspectives and methodologies, mirroring the collaborative nature of real-world drug discovery teams. This novel integration of AI agents with specialized roles represents a significant advancement in the field of computational drug repurposing.
2 Related Works
The concept of drug repurposing has evolved with the advancements in computational tools, leading to a growing body of literature that explores various methodologies. Here, we highlight key developments in the field that align with our multi-agent system approach.
Machine Learning in Drug Repurposing. Machine learning techniques have been applied to drug repurposing, demonstrating significant potential in predicting drug-disease interactions (Issa et al., 2021). Similarly, our AI Agent leverages the DeepPurpose toolkit, which has been extensively validated for its efficiency in drug-target interaction (DTI) predictions (Huang et al., 2020).
Knowledge Graphs for Integrative Analysis. Knowledge graphs have been instrumental in providing a structured way of integrating diverse biological data. For instance, the DRKG, as employed by our Knowledge Graph Agent, integrates data from several sources, including DrugBank (Knox et al., 2024), Hetionet (Himmelstein et al., 2017), and STRING (Szklarczyk et al., 2023), to offer comprehensive insights into possible drug-disease links (Ioannidis et al., 2020). This structured integration facilitates the systematic exploration of potential drug repurposing candidates.
Literature Search for Drug Repurposing. The automation of literature review and data extraction using AI tools, particularly large language models (LLMs), has become an essential component of modern drug discovery (Chakraborty et al., 2023). This trend is relevant in the field of drug repurposing, where efficient processing of vast amounts of scientific literature is crucial. Recent studies have demonstrated that LLM-based search tools can enhance the efficiency and complexity of queries compared to traditional search engines (Spatharioti et al., 2023). In our framework, we leverage this approach by implementing a Search Agent that utilizes search engines as a data source, aligning with current trends in AI-assisted literature review for drug discovery and repurposing.
Multi-Agent Systems in Biomedical Applications. While individual AI applications have shown promise, the integration of these technologies through a multi-agent system is less explored in the field of drug repurposing. However, similar multi-agent frameworks have been implemented in other areas of biomedical research, such as clinical trials (Yue & Fu, 2024). These studies provide a foundation for the application of multi-agent systems in drug repurposing, suggesting that such frameworks can enhance the predictive accuracy and efficiency of computational drug discovery.
3 Methods
Our DrugAgent framework is designed to mimic the collaborative and multidisciplinary nature of drug discovery teams. Each agent in the system is specialized to handle specific aspects of the drug repurposing process, allowing for a more comprehensive and nuanced analysis than traditional single-model approaches.
3.1 Overview of DrugAgent
Our proposed system is a conversational multi-agent framework analogous to a specialized research team focused on drug repurposing. Each agent within this system plays a distinct role, mirroring the specialization seen in pharmaceutical research—some focus on machine learning models, others on search-based analysis, and another is dedicated to knowledge graph exploration.
Figure 1: DrugAgent framework architecture for advanced DTI analysis. This system combines a user-friendly interface with sophisticated internal decision-making processes. It features a central “Coordinator” managing specialized agents: a “Knowledge Graph Agent” accessing biomedical databases (DrugBank, CTD, DGIdb, STITCH), a “Search Agent” utilizing web search engines, and an “AI Agent” employing deep learning models (trained on Davis, Kiba, BindingDB datasets with GNN, CNN, Transformers, etc). The system integrates RDKit and UniProt ID for chemical and protein data processing, culminating in a scoring function that synthesizes multi-source information to generate comprehensive answers for complex drug-target queries.
The system comprises the following key agents:
  1. 1.
    A Coordinator Agent that oversees the specialized agents and integrates their findings.
  2. 2.
    An AI Agent specializing in predicting drug repurposing potential using machine learning models.
  3. 3.
    A Search Agent focusing on analyzing existing literature and data for repurposing opportunities.
  4. 4.
    A Knowledge Graph (KG) Agent dedicated to exploring connections between drugs, diseases, and biological pathways.
To process natural language inputs and generate responses that are coherent and contextually appropriate, each agent utilizes large language models. The system’s reasoning capabilities are enhanced by incorporating methodologies that allow for step-by-step problem-solving and decision-making.
Following the reasoning process, the system is capable of taking actions such as calculating repurposing scores, performing literature-based analyses, and querying knowledge graphs for novel drug-disease associations. By integrating this information through a weighted average approach, the system simulates a highly knowledgeable drug repurposing researcher.
Working in concert, these agents can deliver precise, explainable assessments of drug repurposing potential in response to user inquiries about specific drugs or diseases. This framework’s modular nature allows for easy expansion and refinement of the system’s capabilities in the context of drug repurposing research.
3.2 Agent Roles and Responsibilities
The DrugAgent framework integrates a diverse array of specialized agents, each employing the ReAct (Yao et al., 2022) and LEAST-TO-MOST (Zhou et al., 2022) reasoning methods to meticulously plan their actions. Through the use of advanced search capabilities, access to specialist models, and indexing in databases, these agents can execute a wide range of tasks effectively. Below, we delve into the specific roles and responsibilities assigned to each agent within the system.
3.3 AI Agent
Our approach begins with the AI Agent, which utilizes the MPNN_CNN_BindingDB model from DeepPurpose (Huang et al., 2020) to predict potential drug repurposing opportunities. This model combines Message Passing Neural Networks (MPNN) (Gilmer et al., 2017) for processing molecular structures with Convolutional Neural Networks (CNN) for capturing binding site features. It is trained on the comprehensive BindingDB dataset, which contains a wealth of binding affinity data for DTIs.
The MPNN_CNN_BindingDB model operates as follows:
  1. 1.
    The MPNN component processes the molecular graph of the drug, capturing its structural features.
  2. 2.
    The CNN component analyzes the binding site information of the target protein.
  3. 3.
    These features are then combined and processed through fully connected layers to predict binding affinity or repurposing potential.
The AI Agent’s predictions are continually refined through an iterative training and validation process, using cross-validation techniques to ensure accuracy and robustness.
3.4 Knowledge Graph (KG) Agent
Concurrently, the Knowledge Graph (KG) Agent employs DGIdb (Cannon et al., 2024), DrugBank (Knox et al., 2024), CTD (Davis et al., 2023), and STITCH (Kuhn et al., 2007). From these datasets, we make use of the DTI table and then create the vast drug-gene interaction table. From this, we calculate the number of hops to reach from the drug to the target using the below formula,
$\text{DTI}_{\text{score}}(d,t)=\begin{cases}0&\text{if }d\notin G\text{ or }t\notin G,\\ 0&\text{if }h(d,t)=-1,\\ 1&\text{if }h(d,t)=1,\\ \frac{1}{\ln(1+h(d,t))}&\text{otherwise},\end{cases}$ (1)
where $d$ is a drug, $t$ is a target $G$ is a knowledge graph $h(d,t)$ is a number of hops in the shortest path between $d$ and $t$ in $G$ and $\ln(\cdot)$ is a natural logarithm.
3.5 Search Agent: Information Extraction from Biomedical Literature
Parallel to these processes, the Search Agent leverages large language models to automate the extraction of relevant information from biomedical literature, including new findings published in databases like PubMed. This agent applies natural language processing techniques to extract and annotate data regarding drug efficacy, safety, and novel interactions, which are critical for validating and updating the predictions generated by the other agents.
The search agent’s core functionality can be summarized as follows:
  1. 1.
    Google Search Query: The agent formulates a search query combining the drug name and target name, along with the term “interaction”.
  2. 2.
    Web Scraping: It performs a Google search using this query and scrapes the search results, including titles, links, and snippets.
  3. 3.
    Text Analysis: The agent analyzes the scraped text for the presence of the drug name, target name, and predefined keywords related to interactions and efficacy.
  4. 4.
    Scoring: Based on the presence of these elements, it assigns a score to each search result. The scoring system considers: (1) Presence of both drug and target names; (2) Occurrence of interaction-related keywords; (3) The presence of words indicating strong or significant effects.
  5. 5.
    DTI Score Calculation: Finally, it calculates an overall DTI score by aggregating individual result scores and normalizing the total.
This simplified approach allows for rapid information gathering from publicly available sources. However, it is important to note that this method relies on the quality and relevance of Google search results, and does not directly access or analyze full scientific papers or curated databases. As such, it serves as a preliminary screening tool rather than a comprehensive literature review system.
The DTI score calculation is as follows: Let $R=\{r_{1},r_{2},...,r_{n}\}$ be the set of search results, where $n$ is the number of results (default $n$ is 10). For each result $r_{i}$, we define an individual score function $S(r_{i})$: $S(r_{i})=I(d,t,r_{i})+I(p,r_{i})+I(s,r_{i})$, where
$I(d,t,r_{i})=\begin{cases}1&\text{if drug name }d\text{ and target name }t\text{ are in }r_{i}\\ 0&\text{otherwise},\end{cases}$
$I(p,r_{i})=\begin{cases}1&\text{if any positive keyword is in }r_{i}\\ 0&\text{otherwise},\end{cases}$
and
$I(s,r_{i})=\begin{cases}1&\text{if any strong keyword is in }r_{i}\\ 0&\text{otherwise}.\end{cases}$
The positive keywords are “interacts”, “binds”, “activates”, “inhibits”, and “modulates”. The strong keywords are “strong”, “significant”, “potent”, and “effective”.
The total score $T$ is then calculated as $T=\sum_{i=1}^{n}S(r_{i})$. The maximum possible score $M$ is $M=3n$. Finally, the normalized DTI score $D$ is calculated as:
$D=\begin{cases}\text{round}\left(\frac{T}{M},2\right)&\text{if }M>0\\ 0&\text{if }M=0,\end{cases}$ (2)
where $\text{round}(x,2)$ rounds $x$ to 2 decimal places.
3.6 Calling External Tools
GPT supports calling external tools (e.g., function, database retrieval) to leverage external knowledge and enhance its capability. Specifically, suppose we have a couple of toolkits, GPT’s API can automatically detect which tool to use, which serves as glue to connect large language models to external tools. Our system integrates a variety of external data sources and predictive AI models to support the agents’ functions.
Data Sources
The use of professional datasets is pivotal in ensuring the accuracy and reliability of our agents’ information retrieval capabilities.
  • DrugBank: DrugBank (Knox et al., 2024) stands out as a premier resource, offering detailed drug data, including chemical, pharmacological, and pharmaceutical information, with a focus on comprehensive DTIs. DrugBank is not only a repository of drug information but also serves as an invaluable tool for bioinformatics and cheminformatics research. It provides data for over 13,000 drug entries, including FDA-approved small-molecule drugs, FDA-approved biopharmaceuticals (proteins, peptides, vaccines, and allergens), and nutraceuticals.
  • Comparative Toxicogenomics Database (CTD): The CTD (Davis et al., 2023) is a robust, manually curated database that provides information about chemical–gene/protein interactions, chemical–disease, and gene-disease relationships. It is valuable for understanding how environmental exposures affect human health, integrating data from various species and linking chemicals, genes, diseases, phenotypes, and pathways (Chang et al., 2019; Wu et al., 2022a).
  • Search Tool for Interactions of Chemicals (STITCH): STITCH (Kuhn et al., 2007) is a database of known and predicted interactions between chemicals and proteins. It integrates information from various sources, including experimental data, predictive methods, and text-mining of scientific literature. STITCH is useful for exploring the complex network of interactions between drugs, other chemicals, and proteins.
  • Drug-Gene Interaction Database (DGIdb): DGIdb (Cannon et al., 2024) is a web resource that consolidates disparate data sources describing drug-gene interactions and gene druggability. It provides an intuitive interface for searching drug-gene interactions and potentially druggable genes, making it an essential tool for researchers in fields such as cancer informatics, drug repurposing, and personalized medicine (Chen et al., 2021; Wang et al., 2024; Lu et al., 2024b).
Predictive AI Models
We utilize DeepPurpose (Huang et al., 2020) for the AI Agent. DeepPurpose is a comprehensive and extensible deep learning library for DTI prediction. It integrates multiple state-of-the-art models and datasets, allowing researchers to easily implement and compare various deep learning approaches for drug discovery and repurposing. DeepPurpose facilitates the application of AI in drug development by providing a unified framework for different drug and protein encoding methods.
3.7 Drug Repurposing Agent
The predictions from the AI agent are cross-referenced and enriched with data from the Knowledge Graph Agent and validated against findings from the Search Agent. This integrated workflow allows for a dynamic updating mechanism, wherein feedback from the literature and knowledge graph continually refines the predictions. The entire process is supported by a central data hub that manages data consistency and real-time updates, ensuring that all agents have access to the latest and most accurate data available.
The final prediction score is calculated in this formula:
$S_{merged}=\alpha S_{ml}+\beta S_{search}+(1-\alpha-\beta)S_{kg},$
where $S_{merged}$ is the merged DTI score, $S_{ml}$ is the machine learning DTI score, $S_{search}$ is the search-based DTI score, $S_{kg}$ is the knowledge graph DTI score, $\alpha$ is the weight for the machine learning score, $\beta$ is the weight for the search score, and $(1-\alpha-\beta)$ is the weight for the knowledge graph score. Note that $10<\alpha<$ 1, $10<\beta<$ 1, and $\alpha+\beta<1$.
3.8 Workflow
The workflow of our DrugAgent is designed to leverage the strengths of multiple specialized agents to provide comprehensive and accurate DTI scores. The process is structured in several sequential steps, as described below:
Step 1: Query Initialization and Agent Preparation. The workflow begins with the user input, specifying the drug name, target name, and weighting parameters (alpha and beta). The system initializes four specialized agents: the AI Agent, Search Agent, Knowledge Graph (KG) Agent, and Coordinator Agent. Each agent is configured with specific roles and access to relevant functions and databases.
Step 2: Task Allocation to Specialist Agents. The Coordinator Agent, acting as the central manager, allocates specific tasks to each specialist agent:
  • The AI Agent is tasked with calculating the DTI score using machine learning models.
  • The Search Agent is responsible for analyzing DTI data using search methods and literature analysis.
  • The KG Agent is assigned to analyze DTI data using Knowledge Graph techniques.
Step 3: Independent Agent Processing. Each specialist agent processes its assigned task independently, utilizing its specific methodologies and tools:
  • The AI Agent applies machine learning models to predict the DTI score.
  • The Search Agent conducts literature searches and analyzes the results to derive a DTI score.
  • The KG Agent queries and analyzes the knowledge graph to determine the DTI score.
Step 4: Score Collection and Synthesis. After each agent completes its task, the individual DTI scores are reported back to the Coordinator Agent. The Coordinator synthesizes these scores, applying the provided weighting parameters ($\alpha$ and $\beta$) to merge the individual scores into a final, comprehensive DTI score.
Step 5: Result Integration and Final Output. The Coordinator Agent integrates all the information, including the individual scores from each method and the merged final score. It formats this information into a structured output, providing a comprehensive view of the DTI prediction from multiple perspectives.
Step 6: Delivery of Solution. The final output, which includes the merged DTI score along with the individual scores from each method, is delivered to the user. This comprehensive result provides not only the final prediction but also insights into how different methods contribute to the overall score, enhancing the user’s understanding of the DTI prediction. This structured workflow ensures that our multi-agent DTI prediction system effectively combines multiple analytical approaches, offering a robust and multi-faceted assessment of potential DTIs.
4 Experiment
In this section, we demonstrate the experimental results and case studies. Due to the page limit, the experimental setups, including dataset description, evaluation metrics, and implementation details, are elaborated in the Appendix.
4.1 Quantitative Results: Performance Comparison of DrugAgent and GPT-4
To evaluate the performance of DrugAgent against GPT-4 in predicting PKD scores from the BindingDB database, we compared their predictions using several key statistical metrics. Table 1 summarizes the results for 10 diverse drug-target combinations not used in parameter tuning.
Table 1: Performance comparison between DrugAgent and GPT-4 for different datasets. We report the average results of 5 independent runs and the corresponding standard deviations (in brackets). For each metric, we highlight the best method in bold. We marked the metrics where DrugAgent is significantly better than GPT-4 (pass the t-test, i.e., p-value$<$0.05) using “*”.
Metric DrugAgent GPT-4 w.o. AI Agent w.o. KG Agent w.o. Search Agent
[…]
Runtime ($\downarrow$) $\approx$5.000s $\approx$0.297s - - -
# OpenAI API tokens ($\downarrow$) $\approx$2000-3000 $\approx$100 - - -
cost of tokens ($\downarrow$) $\approx$\$0.006-\$0.027 $\approx$\$0.0014-\$0.0020 - - -
DrugAgent outperformed GPT-4 across all examined metrics. Regarding prediction accuracy, DrugAgent achieved a Mean Squared Error (MSE) of 1.836, significantly lower than GPT-4’s 13.420, indicating superior overall predictive power. The Mean Absolute Percentage Error (MAPE) and Mean Absolute Error (MAE) metrics further confirm this superiority, with DrugAgent achieving values of 0.134 and 1.081, respectively, compared to GPT-4’s 0.320 and 3.350.
DrugAgent demonstrated stronger explanatory power, evidenced by its positive R-squared (R2) value of 0.431, indicating a moderate fit to the data. In contrast, GPT-4’s R2 value of -0.460 suggests a poor fit. The Explained Variance metric reinforces this trend, with DrugAgent achieving a positive value of 0.577, while GPT-4 showed a value of -0.460. These results highlight DrugAgent’s superior ability to capture and explain the variance in the PKD score data.
DrugAgent also excels in prediction consistency. Its Max Error of 2.809 is less than half of GPT-4’s 6.490, indicating more reliable predictions across the dataset. Moreover, the strong positive correlation (0.761) between DrugAgent’s predicted and actual values, compared to GPT-4’s weak positive correlation (0.110), underscores DrugAgent’s effectiveness in capturing the underlying relationships in the data.
These results collectively showcase DrugAgent’s significantly improved predictive capabilities for BindingDB PKD scores compared to GPT-4, which is crucial for accurate binding affinity predictions in drug discovery and molecular interaction studies.
Our ablation study provides valuable insights into the importance of each component in the DrugAgent architecture. Removing the AI Agent resulted in the most severe performance degradation across all metrics, with MSE increasing dramatically to 52.349 and R2 declining to -15.228. This underscores the AI Agent’s critical role in understanding complex patterns in molecular structures and their relationship to binding affinities. The KG Agent also proved essential, as its removal led to significant performance drops, though less severe than the AI Agent. This indicates the KG Agent’s crucial contribution of domain knowledge about chemical structures and known interactions. While the Search Agent had a less dramatic impact on performance, it still contributed to the overall accuracy of the model, particularly in maintaining high correlation and explained variance.
DrugAgent achieves its superior performance with a runtime of 5.000s, which is efficient considering the complexity of its multi-agent architecture. In comparison, GPT-4’s runtime of 0.297s is faster, but at the cost of significantly reduced accuracy.
We also compared the number of OpenAI API tokens used and the associated costs. DrugAgent uses between 2000-3000 tokens per prediction, with an approximate cost of \$0.006-\$0.027, while GPT-4 uses around 100 tokens, costing $0.0014-$0.0020 per prediction. While DrugAgent has higher token usage and cost, its superior performance justifies this increased resource utilization for applications requiring high accuracy.
This comprehensive analysis highlights the synergistic interaction between the AI, KG, and Search Agents in our model architecture for accurate PKD score prediction. The results validate our design choices and provide valuable insights for future improvements and optimizations of DrugAgent. They particularly emphasize the potential for enhancing the integration and capabilities of each component to achieve even more accurate PKD score predictions in the context of the BindingDB database, while considering the balance between performance, computational resources, and cost.
4.2 Case Study
This study presents three case studies analyzing the drug repurposing potential of Topotecan for different proteins: TOP1, SLFN11, and SLC26A4. These cases represent a spectrum from known strong interactions to potentially novel connections, allowing for a comprehensive evaluation of our multi-agent system’s capabilities. In each case, the Coordinator Agent decomposed the complex problem into three subproblems, each handled by a specialized agent. The AI Agent predicted interaction potential based on structural properties, the Search Agent analyzed evidence from scientific literature, and the Knowledge Graph Agent evaluated connections in existing databases.
Case 1 examined the interaction between Topotecan and TOP1, a known strong interaction. The system calculated a final score of 11.51, confirming the established relationship. The high AI Agent score (7.65) and Knowledge Graph Agent score (1.0) aligned with the known mechanism of Topotecan as a TOP1 inhibitor, while the relatively low Search Agent score (0.27) likely reflected the well-established nature of this interaction not requiring extensive new studies.
Case 2 investigated the less understood but potentially relevant interaction between Topotecan and SLFN11. The system yielded a final score of 10.30, suggesting a noteworthy relationship. The high AI Agent score (7.36) indicated structural compatibility, while the moderate Search Agent (0.33) and Knowledge Graph Agent (0.72) scores reflected some existing evidence and established connections, albeit less comprehensive than in Case 1.
Case 3 explored an unlikely interaction between Topotecan and SLC26A4, resulting in a final score of 9.92. Despite a high AI Agent score (7.61) suggesting structural compatibility, the very low Search Agent score (0.00) indicated minimal literature evidence. The moderate Knowledge Graph Agent score (0.72) suggested some indirect connections, highlighting the system’s ability to detect potential novel interactions while also recognizing the need for careful interpretation.
These case studies demonstrated the multi-agent system’s ability to handle diverse scenarios, from known interactions to potential discoveries. The system effectively integrated various data sources and analytical methods, providing interpretable results with detailed reasoning processes. Importantly, it showed the capacity to recognize its own limitations, particularly when predictions conflicted with existing evidence.
The system offered practical insights for drug repurposing research across different levels of prior knowledge. In Case 1, it confirmed a well-established interaction. In Case 2, it suggested a potentially relevant interaction that warrants further investigation. In Case 3, it identified a possible novel interaction while appropriately flagging the low literature evidence, thus highlighting an area requiring careful experimental validation.
This comprehensive approach not only validates known interactions but also suggests potential new avenues for research. By providing nuanced analyses that consider multiple factors, the system demonstrates its value in accelerating and refining the drug repurposing process. Future work could focus on experimental validation of the system’s predictions, particularly for cases like Topotecan-SLFN11 where moderate evidence suggests a promising avenue for investigation.
5 Discussion
Our study presents a novel multi-agent system for drug repurposing that integrates machine learning, knowledge graphs, and literature search. This approach offers more robust predictions by leveraging diverse data sources and analytical methods. The system’s strength lies in its collaborative approach, which combines each agent’s specialized capabilities to evaluate complex DTIs comprehensively. The weighted integration method allows for flexible adjustment of different prediction methods, enhancing overall accuracy. However, limitations exist. The system still relies on human expertise for initial setup, limiting its scalability. It also lacks autonomous knowledge updating capabilities to keep pace with rapidly evolving pharmacological research. Furthermore, the current model does not adequately account for individual patient characteristics or drug combination effects.
Future research should focus on integrating autonomous knowledge updating mechanisms to keep the system current with the latest pharmacological research. It should also enhance the system’s ability to predict drug efficacy and handle complex drug combinations. Expanding the system’s applicability to a broader range of pharmaceutical tasks by incorporating existing models is crucial. This includes adapting the framework for Drug Response Prediction using models like drGAT (Inoue et al., 2024b), extending to DTI prediction tasks, leveraging the system’s existing knowledge graph and AI capabilities, and developing a more flexible prompting system to allow easy adaptation to various pharmaceutical tasks without major architectural changes, such as drug synergy prediction (Huang et al., 2022), drug property prediction (Xu et al., 2024), adverse drug reaction prediction (Chen et al., 2024a) or drug design (Fu et al., 2021a; 2022a).
Automating the preprocessing of complex data types, such as scRNA-seq data, is another important area for improvement. This involves implementing multiple preprocessing functions including Imputation (Inoue et al., 2024a; Inoue, 2024), Quality Control (McCarthy et al., 2017; Lu et al., 2023), and batch effect correction (Li et al., 2020; Fu et al., 2024; Haghverdi et al., 2018). Developing an optimization framework to automatically select and apply the most appropriate preprocessing methods for given datasets and integrating these preprocessing capabilities seamlessly into the existing multi-agent system are essential steps.
Validating the expanded system’s predictions through experimental studies and clinical trials to ensure real-world applicability across multiple tasks and data types (Lu et al., 2024a; Fu et al., 2023) is crucial for future development.
In conclusion, our DrugAgent shows promise in accelerating AI-driven drug discovery. Continued development addressing both computational and pharmacological challenges could lead to more efficient and cost-effective drug discovery processes(Zhang et al., 2021). Future work should focus on validating the system in real-world drug discovery projects (Fu et al., 2021b) and evaluating its performance with larger, more diverse datasets.
Appendix A Implementation
A.1 Pattern Selection and Weight Optimization for Score Integration
To determine the optimal integration method for our agent scores, we explored four different mathematical patterns and employed a constrained optimization method. We utilized a dataset comprising 3,332 drug-target pairs, each containing scores from our three specialized agents (AI, Knowledge Graph, and Search) along with corresponding ground truth interaction scores (PKD values) from the BindingDB dataset. We considered the following four patterns for score integration:
$f_{1}(\alpha,\beta,\gamma,A,B,C)$ $=\alpha A+\beta B+\gamma C,$ (3)
$f_{2}(\alpha,\beta,\gamma,A,B,C)$ $=\alpha A+(\beta B\cdot\gamma C),$
$f_{3}(\alpha,\beta,\gamma,A,B,C)$ $=(\alpha A\cdot\beta B)+\gamma C,$
$f_{4}(\alpha,\beta,\gamma,A,B,C)$ $=(\alpha A\cdot\gamma C)+\beta B,$
where $A$, $B$, and $C$ represent the scores from the AI, Knowledge Graph, and Search agents respectively, and $\alpha$, $\beta$, and $\gamma$ are the weights we are optimizing. For each pattern, we formulated the weight optimization as a constrained minimization problem:
$\underset{\alpha,\beta,\gamma}{\text{minimize}}$ $\|Y-f(\alpha,\beta,\gamma,A,B,C)\|_{2}$ (4)
subject to $\alpha,\beta,\gamma\geq 0,$
where $Y$ is the vector of ground truth PKD values and $i\in{1,2,3,4}$ corresponds to the pattern type. The optimization was performed using Sequential Least Squares Programming (SLSQP) with non-negative constraints for each pattern. After comparing the results, we found that the linear combination pattern $f_{1}$ yielded the best performance:
$f_{1}(\alpha,\beta,\gamma,A,B,C)=\alpha A+\beta B+\gamma C$ (5)
The initial optimized weights for this pattern were:
$\alpha$ $=1.24683589$ (6)
$\beta$ $=2.23513134$
$\gamma$ $=3.22163745\times 10^{-16}$
After obtaining these initial results, we rounded and adjusted the coefficients for practical implementation and to account for potential overfitting to our specific dataset. The final weights used in our merged DTI score calculation are:
$S_{\text{merged}}=1.2S_{\text{AI}}+2.2S_{\text{KG}}+0.5S_{\text{Search}}$ (7)
The selection of the linear combination pattern ($f_{1}$) and the subsequent weight adjustments reflect several important considerations:
The linear pattern provided the best fit to our data while maintaining simplicity and interpretability. The Knowledge Graph (KG) agent retains the highest weight, confirming its significant contribution to the final prediction. The AI agent continues to play a substantial role, with a weight slightly lower than the KG agent. While the initial optimization suggested a negligible role for the Search agent, we maintained its contribution at a non-trivial level (0.5) in the final model. This adjustment reflects our belief in the potential value of diverse information sources, even if not prominently represented in our current dataset.
It is worth noting that the sum of these weights (3.9) is intentionally not normalized to 1. This allows for a more flexible scaling of the final score, which can be beneficial in certain applications or when comparing across different datasets. This approach, combining pattern selection, data-driven weight optimization, and expert adjustment, ensures that our final predictions leverage the strengths of each agent while maintaining robustness and generalizability. The results suggest that a linear combination of structured knowledge from the KG agent, pattern recognition capabilities of the AI agent, and supplementary information from the Search agent contributes to accurate PKD value prediction, with the KG and AI agents playing particularly crucial roles.
A.2 Implementation Details
In this section, we provide detailed descriptions of the implementation processes to enhance the reproducibility of our study.
Role Assignment to Agents
Each agent within our multi-agent framework is designated a specific role, which is integrated directly into the LLM’s system prompt for clarity and focus. For instance, the role of the AI Agent is defined as follows:
”””Specialized AI Agent for calculating DTI scores using machine learning models. Use the get_AI_score function to obtain the DTI score. Output the score in the following format: { ”AI_score”: 1.0,}”””
This role definition is crucial as it guides the LLM to prioritize responses based on the assigned expert domain, leveraging the model’s inherent capability to focus more acutely on instructed tasks than on general information.
Defining External Tools
External tools are defined in a structured format to facilitate their integration and usage within the LLM environment. These definitions are crafted in Python functions, specifying the function name, parameters, and return types. Key examples include:
  1. 1.
    AI Agent: Utilizes machine learning models for DTI scoring.
  2. 2.
    Search Agent: Performs web-based information retrieval to gather relevant DTI data.
  3. 3.
    KG Agent: Leverages a knowledge graph (KG) for graph-based DTI scoring.
This structured approach allows for the direct execution of function calls within the system, providing detailed responses, including the function name and arguments. These responses enable the retrieval of results in a structured manner. (See Appendix A for detailed implementations of these agents)
Enhanced Score Integration
To improve the model’s prediction capabilities, we incorporate a weighted integration method within the Coordinator Agent. This method, known as score merging, aids in synthesizing the outputs from different agents into a comprehensive DTI prediction. The integration is performed using predefined weights ($\alpha$, $\beta$, and $\gamma$) to balance the contributions of each prediction method:
$merged\_dti\_score=\alpha*AI\_score+\beta*KG\_score+\gamma*search\_score$
Here, AI_score, search_score, and KG_score represent the DTI scores from the AI Agent, Search Agent, and KG Agent, respectively. The weights $\alpha$, $\beta$ $\gamma$ are adjustable hyperparameters that determine the relative importance of each score in the final prediction. This approach enhances the accuracy of the model’s outputs and its ability to leverage diverse prediction methods for more robust drug repurposing predictions. This parameter and the formula were defined systematically. (See A.1 in detail.)
These implementation strategies collectively ensure that each component of our multi-agent system operates effectively and that the integration between different agents and external tools is seamless, fostering an environment conducive to robust, reproducible research in drug repurposing prediction.
Software and Hardware Configuration
Our experimental framework was implemented on a Mac computer equipped with an Apple M1 chip and 16GB unified memory, utilizing the built-in GPU cores. We used Python 3.10 for scripting, PyAutoGen 0.2.31 (Wu et al., 2023), DeepPurpose 0.1.5 (Huang et al., 2020), and RDKit 2023.9.6 (Landrum et al., 2024). For each experiment, we used the same seed to ensure reproducibility across different Mac models.
A.3 Detailed External Tool Definitions
This appendix provides a comprehensive overview of the implementation details for our three key agents: AI Agent, Search Agent, and KG Agent. Each agent plays a crucial role in our multi-agent system for drug repurposing prediction.
A.3.1 AI Agent Implementation
The AI Agent utilizes machine learning models to predict DTIs. Its core function, get_AI_score, takes a drug name and a target name as input and returns a float value representing the predicted interaction score.
# AI Agent
def get_ml_dti_score(name: str, target_name: str) -> float:
target_sequence = get_target_sequence(target_name)
net = models.model_pretrained(model="MPNN_CNN_BindingDB")
X_repurpose, drug_name, drug_cid = load_broad_repurposing_hub(
SAVE_PATH
idx = drug_name == name
if not any(idx):
print(f"Logging: Drug ’{name}’ not found.")
return None
res = models.virtual_screening(
X_repurpose[idx], [target_sequence], net,
drug_name[idx], [target_name]
return res[0]
This implementation uses a pre-trained MPNN_CNN model from the BindingDB dataset. It first retrieves the target protein sequence and loads the drug data. If the specified drug is found, it performs virtual screening to predict the interaction score.
A.3.2 Search Agent Implementation
The Search Agent leverages web-based information to gather relevant data about DTIs. It consists of several functions that work together to perform a Google search, parse the results, and calculate a DTI score based on the search findings.
# Search Agent
def google_search(query: str, num_results: int = 10) -> List[Dict[str, str]]:
# ... [implementation details]
def _parse_search_results(soup: BeautifulSoup) -> List[Dict[str, str]]:
# ... [implementation details]
def calculate_dti_score(search_results: List[Dict[str, str]],
drug_name: str, target_name: str) -> float:
# ... [implementation details]
def _calculate_individual_score(result: Dict[str, str], drug_name: str,
target_name: str, positive_keywords: List[str],
strong_keywords: List[str]) -> int:
# ... [implementation details]
def analyze_dti(name: str, target_name: str) -> float:
search_results = google_search(f"{name} {target_name} interaction")
dti_score = calculate_dti_score(search_results, name, target_name)
return dti_score
The main function, analyze_dti, orchestrates the search process and score calculation. It uses a keyword-based scoring system to evaluate the relevance and strength of the interaction based on search results.
A.3.3 KG Agent Implementation
The KG Agent utilizes a knowledge graph to derive DTI scores based on the structural relationships between drugs and targets in the graph.
# KG Agent
def calculate_dti_score(kg, drug, target):
if drug not in kg.graph or target not in kg.graph:
return 0 # Return 0 if the drug or target is not in the knowledge graph
hops = kg.shortest_path(drug, target)
if hops == -1:
return 0 # No relationship
elif hops == 1:
return 1 # Direct connection
else:
return 1 / (np.log1p(hops)) # Logarithm-based score
def load_kg(file_path):
with open(file_path, "rb") as f:
kg = pickle.load(f)
return kg
def get_kg_dti_score(name: str, target_name: str) -> float:
kg = load_kg("../data/knowledge_graph.pkl")
score = calculate_dti_score(kg, name, target_name)
return score
The KG Agent loads a pre-constructed knowledge graph and calculates the DTI score based on the shortest path between the drug and target nodes in the graph. A direct connection yields the highest score, while more distant connections result in lower scores, calculated using a logarithmic scale.
These detailed implementations demonstrate how each agent contributes unique insights to the overall DTI prediction task. The AI Agent provides predictions based on learned patterns from large datasets, the Search Agent incorporates up-to-date information from web sources, and the KG Agent leverages structured knowledge representations. By combining these diverse approaches, our system aims to produce more robust and comprehensive drug repurposing predictions.
Appendix B Experimental Setup
B.1 Dataset and Evaluation Metrics
In our study, we utilized the Kd (dissociation constant) data from the BindingDB database (Liu et al., 2007) as our experimental dataset. BindingDB is a public repository of measured binding affinities, primarily focusing on interactions between proteins considered as drug targets and small, drug-like molecules.
The Kd dataset comprises 52,284 DTI pairs, involving 10,665 unique drug-like compounds and 1,413 distinct protein targets. Kd values represent the dissociation constant, which quantifies the propensity of a larger complex to separate (dissociate) into smaller components. A lower Kd value indicates a higher binding affinity between the drug and the target protein.
Our regression task involved predicting PKD (negative logarithm of the dissociation constant Kd) values from the BindingDB database, given the target protein’s amino acid sequence and the drug compound’s SMILES string (Simplified Molecular-Input Line-Entry System, a line notation for encoding molecular structures) (Weininger, 1988). This task is crucial for understanding Drug-Target Interactions (DTIs) and has significant implications for drug discovery and development processes (Huang et al., 2022).
To comprehensively evaluate our model’s performance, we employed a suite of seven complementary metrics: Mean Squared Error (MSE), Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), R-squared (R2) Score, Explained Variance, Maximum Error, and Correlation. These metrics collectively assess various aspects of our predictions: MSE and MAE provide measures of the average prediction error, with MSE being more sensitive to large errors due to its quadratic nature. MAPE offers insight into the relative size of prediction errors. The R2 score and Explained Variance evaluate the model’s capacity to capture the underlying variance in the data. Maximum Error highlights the worst-case prediction scenario, crucial for understanding the model’s limitations. Lastly, Correlation assesses the strength and direction of the relationship between predicted and actual PKD values.
This comprehensive set of metrics allows us to thoroughly assess the accuracy of our predictions, the model’s capacity to capture underlying patterns in DTIs, and its consistency across different scenarios. Such a multi-faceted evaluation is essential for validating the model’s performance and identifying areas for potential improvement in the context of PKD score prediction for drug-target interactions. Due to space limitation, implementation details are provided in Appendix (Section A.2).
B.2 Procedure
Each agent in our DrugAgent system was tasked with specific roles, as outlined in the Methods section. The AI Agent applied machine learning models to calculate the DTI score, the Search Agent analyzed literature data to derive a DTI score based on published research, and the KG Agent evaluated DTIs using graph-based techniques. The Coordinator Agent then synthesized these findings into a comprehensive DTI prediction. We conducted experiments to assess the accuracy of the merged DTI scores and the consistency of predictions across different methods.